# Gauging Non-Invertible Symmetries: Topological Interfaces and Generalized Orbifold Groupoid in 2d QFT

### 22 Dec 2023 v2

@@ 2.6 Simple Dimensional Constraints for Generalized Gauging
This relation (together with (2.30)) will become handy to identity the dual algebra object $A^{*}$ in explicit examples. For sequential gauging ${\mathcal{C}}\to{\mathcal{C}}^{\prime}\to{\mathcal{C}}^{\prime\prime}$ implemented by haploid algebras $A$ and $A^{\prime}$ successively with corresponding simple topological interfaces ${\cal I}$ and ${\cal I}^{\prime}$, if the topological interface between ${\mathcal{C}}$ and ${\mathcal{C}}^{\prime\prime}$ resulting from the fusion ${\cal I}^{\prime\prime}={\cal I}\otimes{\cal I}^{\prime}$ remains simple, [^20] This is not the case if ${\mathcal{C}}^{\prime\prime}\cong{\mathcal{C}}$ and ${\cal I}^{\prime}\cong\overline{{\cal I}}$ which represents the topological interface that undoes the $A$ gauging by $A^{*}$, and (2.41) does not apply. the one-step gauging is simply implemented by the haploid algebra $A^{\prime\prime}={\cal I}^{\prime\prime}\otimes\overline{{\cal I}^{\prime\prime}}$, which implies that, in particular,
@@ 3 Generalized Gauging in ${\rm\bf Rep}(H_{8})$ and ${\rm\bf Rep}(D_{8})$
=${}_{g,h}={1\over 2}\chi(g,h)^{-1}\,,\quad[F_{{\mathcal{V}}g{\mathcal{V}}}^{h}]_{{\mathcal{V}},{\mathcal{V}}}=[F_{g{\mathcal{V}}h}^{\mathcal{V}}]_{{\mathcal{V}},{\mathcal{V}}}=\chi(g,h)\,,$ (3.5)
@@
and ${\gamma}\in H^{3}(D_{8},U(1))\cong{\mathbb{Z}}_{4}\times{\mathbb{Z}}_{2}^{2}$ is an order-two 3-cocycle taking values in {±$\{\pm 1\}\subset U(1)$ as follows: ${\gamma}(s^{m_{1}}r^{n_{1}},s^{m_{2}}r^{n_{2}},s^{m_{3}}r^{n_{3}})=1$ unless $m_{1}$ is odd and we are in one of the following three cases: $n_{2}=1$ and $s^{m_{3}}r^{n_{3}}\in\{r^{2},sr^{3}\}$; $n_{2}=2$ and $s^{m_{3}}r^{n_{3}}\in\{r,r^{2},sr^{3},s\}$; $n_{2}=3$ and $s^{m_{3}}r^{n_{3}}\in\{r^{3},sr^{2}\}$. [^23] We have explicitly checked that this is desired cocycle using GAP [100] and SageMath [101] and applying the universal coefficient theorem for group cohomology.
@@ 3.1 NIM-reps for ${\rm\bf TY}({\mathbb{Z}}_{2}^{2})$ Fusion Ring
${\mathcal{L}}_{i}\otimes A\otimes\overline{{\mathcal{L}}}_{j}={\mathcal{L}}_{i}\otimes({\cal I}_{a}\otimes\overline{{\cal I}}_{a})\otimes\overline{{\mathcal{L}}}_{j}\,.$ (3.8)
Projecting onto the ground state on the circle (e.g. taking the cylinder to be thin), we have
$M_{ij}\equiv\dim{\rm Hom}({\mathcal{L}}_{i}\otimes A,{\mathcal{L}}_{j})=\mathop{\rm tr}\nolimits(P_{i}\cdot P_{j}^{t})\,,$ (3.9)
where the trace is taken over the basis of simple objects $\{{\cal I}_{a}\}$ in the putative ${\mathcal{C}}$-module category ${\mathcal{C}}_{A}$. The matrix $M_{ij}$ is easily determined from the fusion ring and the candidate $A$. We then perform matrix factorization to find possible NIM-rep matrices $P_{i}$ satisfying (2.21) of dimension $r$ subject to the dimensional constraints in Section 2.6, which also further restricts $A$. Each NIM-rep =${\mathcal{M}}=\{P_{i}\}$ derived this way has a corresponding dimension vector $d_{\mathcal{M}}$ from the Frobenius-Perron eigenvector (see Section 2.6). Two NIM-reps =${\mathcal{M}}=\{P_{i}\}$ and =${\mathcal{M}}^{\prime}=\{P^{\prime}_{i}\}$ of the same dimension $r$ are equivalent if they are related by an $S_{r}$ permutation of $\{{\cal I}_{a}\}$ (in particular $d_{\mathcal{M}}=d_{{\mathcal{M}}^{\prime}}$ up to permutation). The above procedure produces a collection of irreducible and inequivalent NIM-reps which are candidates for full-fledged module categories for the fusion categories that share the input fusion ring. For each of these NIM-irreps, the corresponding algebra objects can be read-off from the diagonal entries of $P_{i}$ as in (2.32).
@@
${\rm NR}_{1}=$$\left\{\left(\begin{array}[]{rrrrr}1&0&0&0&0\\ 0&1&0&0&0\\ 0&0&1&0&0\\ 0&0&0&1&0\\ 0&0&0&0&1\end{array}\right)\,,\left(\begin{array}[]{rrrrr}0&1&0&0&0\\ 1&0&0&0&0\\ 0&0&0&1&0\\ 0&0&1&0&0\\ 0&0&0&0&1\end{array}\right)\,,\left(\begin{array}[]{rrrrr}0&0&1&0&0\\ 0&0&0&1&0\\ 1&0&0&0&0\\ 0&1&0&0&0\\ 0&0&0&0&1\end{array}\right)\,,\right.$ (3.11)
@@
${\rm NR}_{2}=$$\left\{\left(\begin{array}[]{rrrr}1&0&0&0\\ 0&1&0&0\\ 0&0&1&0\\ 0&0&0&1\end{array}\right)\,,\left(\begin{array}[]{rrrr}0&1&0&0\\ 1&0&0&0\\ 0&0&1&0\\ 0&0&0&1\end{array}\right)\,,\left(\begin{array}[]{rrrr}0&1&0&0\\ 1&0&0&0\\ 0&0&0&1\\ 0&0&1&0\end{array}\right)\,,\left(\begin{array}[]{rrrr}1&0&0&0\\ 0&1&0&0\\ 0&0&0&1\\ 0&0&1&0\end{array}\right)\,,\right.$ (3.12)
@@
${\rm NR}_{3}=$$\left\{\left(\begin{array}[]{rrrr}1&0&0&0\\ 0&1&0&0\\ 0&0&1&0\\ 0&0&0&1\end{array}\right)\,,\left(\begin{array}[]{rrrr}0&1&0&0\\ 1&0&0&0\\ 0&0&0&1\\ 0&0&1&0\end{array}\right)\,,\left(\begin{array}[]{rrrr}0&1&0&0\\ 1&0&0&0\\ 0&0&0&1\\ 0&0&1&0\end{array}\right)\,,\left(\begin{array}[]{rrrr}1&0&0&0\\ 0&1&0&0\\ 0&0&1&0\\ 0&0&0&1\end{array}\right)\,,\right.$ (3.13)
@@ 3.2 Algebra objects in ${\rm\bf Rep}(H_{8})$
$A=A^{\rm I}:\quad$ $m_{11}^{1}=m_{1,\eta\eta^{\prime}}^{\eta\eta^{\prime}}=m_{\eta\eta^{\prime},1}^{\eta\eta^{\prime}}=m_{\eta\eta^{\prime},\eta\eta^{\prime}}^{1}=m_{1{\mathcal{V}}}^{{\mathcal{V}}}=m_{{\mathcal{V}}1}^{{\mathcal{V}}}=m_{{\mathcal{V}}{\mathcal{V}}}^{1}$ (3.20)
@@
$A=A^{\rm II}:\quad$ $m_{g,h}^{gh}=\frac{1}{2\sqrt{2}}\begin{pmatrix}1&1&1&1\\ 1&1&-i&i\\ 1&i&1&-i\\ 1&-i&i&1\end{pmatrix}\,,m_{1{\mathcal{V}}}^{{\mathcal{V}}}=m_{{\mathcal{V}}1}^{{\mathcal{V}}}=m_{{\mathcal{V}}{\mathcal{V}}}^{1}=\frac{1}{2\sqrt{2}}\begin{pmatrix}1&0\\ 0&1\end{pmatrix}\,,$ (3.21)
@@
$L\subset D_{8}$ Algebra object $A$ Module Category ${\mathcal{C}}_{A}$ Dual Fusion Category ${}_{A}{\mathcal{C}}_{A}$
@@ 3.3 Module Categories and the Orbifold Groupoid for ${\rm\bf Rep}(H_{8})$
The rank of each indecomposable module category ${\mathcal{M}}$ follows from that of the corresponding NIM-irrep, and in the second column of Table 1 we list the simple objects in ${\mathcal{M}}$ as simple right $A$-modules (written as non-simple TDLs in ${\mathcal{C}}$) using the equivalence ${\mathcal{M}}\cong{\mathcal{C}}_{A}$ [68]. Note that all simple $A$-modules can be determined from (submodules of) induced $A$-modules (namely $A\otimes{\mathcal{L}}_{i}$ for simple TDLs ${\mathcal{L}}_{i}\in{\mathcal{C}}$) in a standard procedure [64], by taking into account possible nontrivial endomorphisms of the induced module that dictates how such an induced module splits into simple $A$-modules. [^24] This is the direct generalization of the usual fixed point resolution in ordinary orbifold [104] for both local operators and defects. Here this splitting happens for the two rank 4 module categories, where the induced module $A\otimes{\mathcal{V}}$ decomposes into two simple $A$-modules that are differentiated by the subscript on ${\mathcal{V}}$, and the corresponding NIM-irreps are given in (3.12) and (3.13). For the algebra object $A=\mathds{1}\oplus\eta\oplus\eta^{\prime}\oplus\eta\eta^{\prime}$ there are two inequivalent multiplication morphisms that differ by the discrete torsion $H^{2}({\mathbb{Z}}_{2}^{2},U(1))={\mathbb{Z}}_{2}$ (differentiated by the subscript $\star$ in Table 1). One of them is Morita equivalent to trivial gauging where the multiplication morphism is the trivial canonical one for $A={\mathcal{V}}\otimes{\mathcal{V}}$, and the induced module $A\otimes{\mathcal{V}}$ splits into four simple $A$-modules. The resulting module category is the regular module category that is universal to any fusion category with the corresponding NIM-irrep (3.11). In the other case, this splitting is forbidden due to the nontrivial multiplication morphism, and the resulting module category is rank 2 and corresponds to the NIM-irrep (3.15). Finally the two non-invertible gaugings where ${\mathcal{V}}\in A$ produce a rank 2 module category and a rank 1 module category with NIM-irreps (3.14) and (3.16), respectively. The rank 1 module category ${\rm\bf Vec}$ with the unique maximal algebra object realizes the unique fiber functor for ${\rm\bf Rep}(H_{8})$ [90].
@@
where the ${\rm\bf Rep}(H_{8})$-module category (^${\mathcal{M}}(\hat{\mathbb{Z}}_{2},1)$ is equivalent to the opposite of the ${\rm\bf Vec}_{D_{8}}^{\gamma}$-module category ${\mathcal{M}}({\mathbb{Z}}_{2}^{s},1)$ as an invertible bimodule category over $({\rm\bf Rep}(H_{8}),{\rm\bf Vec}_{D_{8}}^{\gamma})$. All ${\mathbb{Z}}_{2}$ symmetries in ${\rm\bf Rep}(H_{8})$ are clearly gaugeable, however the special one ^$\hat{\mathbb{Z}}_{2}$ needed here must be the subgroup generated by the invertible TDL $\eta\eta^{\prime}$ which is distinguished by having trivial F-symbols (i.e. 1 if the fusion channel exists) with all TDLs in ${\rm\bf Rep}(H_{8})$. This ensures that the duality defect ${\mathcal{V}}$ under this ^$\hat{\mathbb{Z}}_{2}$ gauging can split consistently into a pair of conjugate invertible TDLs that correspond to generators of the ${\mathbb{Z}}_{4}$ subgroup of dual $D_{8}$ symmetry. We thus identify ${\rm\bf Vec}_{D_{8}}^{\gamma}$ as the dual category for gauging $A=\mathds{1}\oplus\eta\eta^{\prime}$ in Table 1.
@@
as is found in [106]. The full groupoid ${\rm\bf BrPic}$ consists of $328\times 2^{2}=$ 32 invertible bimodule categories (Morita equivalence relations) whose composition laws are all commutative.
@@ 3.4 Algebra objects in ${\rm\bf Rep}(D_{8})$
$L\subset D_{8}$ Algebra object $A$ Module Category ${\mathcal{C}}_{A}$ Dual Fusion Category ${}_{A}{\mathcal{C}}_{A}$
@@
$\langle sr\rangle$ $\mathds{1}\oplus\eta\eta^{\prime}\oplus{\mathcal{V}}$ ${\rm NR}_{4}$: $\{A,\eta\oplus\eta^{\prime}\oplus{\mathcal{V}}_{2}\}$ ${\rm\bf Rep}(D_{8})$
$\langle s\rangle$ $\mathds{1}\oplus\eta^{\prime}\oplus{\mathcal{V}}$ ${\rm NR}_{4}$: $\{A,\eta\oplus\eta\eta^{\prime}\oplus{\mathcal{V}}_{2}\}$ ${\rm\bf Rep}(D_{8})$
@@
Table 2: Algebra objects $A$ in ${\rm\bf Rep}(D_{8})$ grouped according to the Morita equivalence classes together with the corresponding module categories ${\mathcal{C}}_{A}$ (simple objects therein and the corresponding NIM-rep from Section 3.1) and dual fusion categories ${}_{A}{\mathcal{C}}_{A}$. The subscripts on the algebra object $A$ label different multiplication morphisms. The first column labels the module categories using group-theoretical data in $D_{8}$ according to (2.28) for $(H,\psi)=(D_{8},1)$.

### 25 Mar 2026 v3

@@ 2.6 Simple Dimensional Constraints for Generalized Gauging
[…]
@@ 3 Generalized Gauging in ${\rm\bf Rep}(H_{8})$ and ${\rm\bf Rep}(D_{8})$
${}_{g,h}={1\over 2}\chi(g,h)^{-1}\,,\quad[F_{{\mathcal{V}}g{\mathcal{V}}}^{h}]_{{\mathcal{V}},{\mathcal{V}}}=[F_{g{\mathcal{V}}h}^{\mathcal{V}}]_{{\mathcal{V}},{\mathcal{V}}}=\chi(g,h)\,$, (3.5)
@@
and ${\gamma}\in H^{3}(D_{8},U(1))\cong{\mathbb{Z}}_{4}\times{\mathbb{Z}}_{2}^{2}$ is an order-two 3-cocycle taking values in $\{\pm 1\}\subset U(1)$ as follows: ${\gamma}(s^{m_{1}}r^{n_{1}},s^{m_{2}}r^{n_{2}},s^{m_{3}}r^{n_{3}})=1$ unless $m_{1}$ is odd and we are in one of the following three cases: $n_{2}=1$ and $s^{m_{3}}r^{n_{3}}\in\{r^{2},sr^{3}\}$; $n_{2}=2$ and $s^{m_{3}}r^{n_{3}}\in\{r,r^{2},sr^{3},s\}$; $n_{2}=3$ and $s^{m_{3}}r^{n_{3}}\in\{r^{3},sr^{2}\}$. [^23] We have explicitly checked that this is desired cocycle using GAP [100] and SageMath [101] and applying the universal coefficient theorem for group cohomology.
@@ 3.1 NIM-reps for ${\rm\bf TY}({\mathbb{Z}}_{2}^{2})$ Fusion Ring
${\mathcal{L}}_{i}\otimes A\otimes\overline{{\mathcal{L}}}_{j}={\mathcal{L}}_{i}\otimes({\cal I}_{a}\otimes\overline{{\cal I}}_{a})\otimes\overline{{\mathcal{L}}}_{j}\,$, (3.8)
where it’s important to remember that the algebra $A$ may depend on the choice of ${\cal I}_{a}$ (however, the Morita equivalence class of $A$ is fixed). Projecting onto the ground state on the circle (e.g. taking the cylinder to be thin), we have
$M_{ij}\equiv\dim{\rm Hom}({\mathcal{L}}_{i}\otimes A,{\mathcal{L}}_{j})=\sum_{b}P_{ia}{}^{b}P_{jb}{}^{a}\,$, (3.9)
where the matrix $M_{ij}$ is easily determined from the fusion ring and the candidate AA. [^24] Using the Frobenius recipricocity of the fusion ring one can show such an $M_{ij}$ exists for any NIM-rep regardless of whether or not it corresponds to a module category. Therefore, the algorithm detailed here will find all NIM-reps of a given fusion ring. We then perform matrix factorization to find possible NIM-rep matrices $P_{i}$ satisfying (2.21) of dimension $r$ subject to the dimensional constraints in Section 2.6, which also further restricts $A$. Each NIM-rep ${\mathcal{M}}=\{P_{i}\}$ derived this way has a corresponding dimension vector $d_{\mathcal{M}}$ from the Frobenius-Perron eigenvector (see Section 2.6). Two NIM-reps ${\mathcal{M}}=\{P_{i}\}$ and ${\mathcal{M}}^{\prime}=\{P^{\prime}_{i}\}$ of the same dimension $r$ are equivalent if they are related by an $S_{r}$ permutation of $\{{\cal I}_{a}\}$ (in particular $d_{\mathcal{M}}=d_{{\mathcal{M}}^{\prime}}$ up to permutation). The above procedure produces a collection of irreducible and inequivalent NIM-reps which are candidates for full-fledged module categories for the fusion categories that share the input fusion ring. For each of these NIM-irreps, the corresponding algebra objects can be read-off from the diagonal entries of $P_{i}$ as in (2.32).
@@
$\left.\left(\begin{array}[]{rrrrr}0&0&0&1&0\\ 0&0&1&0&0\\ 0&1&0&0&0\\ 1&0&0&0&0\\ 0&0&0&0&1\end{array}\right)\,,\left(\begin{array}[]{rrrrr}0&0&0&0&1\\ 0&0&0&0&1\\ 0&0&0&0&1\\ 0&0&0&0&1\\ 1&1&1&1&0\end{array}\right)\right\}\,,\quad A_{1}=\mathds{1}\,,\mathds{1}\oplus\eta\oplus\eta^{\prime}\oplus\eta\eta^{\prime}\,$,
@@
[…]
@@
[…]
@@ 3.2 Algebra objects in ${\rm\bf Rep}(H_{8})$
$A=A^{\rm I}:$ $m_{11}^{1}=m_{1,\eta\eta^{\prime}}^{\eta\eta^{\prime}}=m_{\eta\eta^{\prime},1}^{\eta\eta^{\prime}}=m_{\eta\eta^{\prime},\eta\eta^{\prime}}^{1}=m_{1{\mathcal{V}}}^{{\mathcal{V}}}=m_{{\mathcal{V}}1}^{{\mathcal{V}}}=m_{{\mathcal{V}}{\mathcal{V}}}^{1}$ (3.20)
@@
$A=A^{\rm II}:$ $m_{g,h}^{gh}=\frac{1}{2\sqrt{2}}\begin{pmatrix}1&1&1&1\\ 1&1&-i&i\\ 1&i&1&-i\\ 1&-i&i&1\end{pmatrix}\,,m_{1{\mathcal{V}}}^{{\mathcal{V}}}=m_{{\mathcal{V}}1}^{{\mathcal{V}}}=m_{{\mathcal{V}}{\mathcal{V}}}^{1}=\frac{1}{2\sqrt{2}}\begin{pmatrix}1&0\\ 0&1\end{pmatrix}\,$, (3.21)
@@
$L\subset D_{8}$ Algebra object AA Module Category ${\mathcal{C}}_{A}$ Dual Fusion Category $_{A}{\mathcal{C}}_{A}$
@@ 3.3 Module Categories and the Orbifold Groupoid for ${\rm\bf Rep}(H_{8})$
The rank of each indecomposable module category $\mathcal{M}$} follows from that of the corresponding NIM-irrep, and in the third column of Table 1 we list the simple objects in $\mathcal{M}$} as simple right $A$-modules (written as non-simple TDLs in $\mathcal{C}$}) using the equivalence ${\mathcal{M}}\cong{\mathcal{C}}_{A}$ [68]. Note that all simple $A$-modules can be determined from (submodules of) induced $A$-modules (namely $A\otimes{\mathcal{L}}_{i}$ for simple TDLs ${\mathcal{L}}_{i}\in{\mathcal{C}}$) in a standard procedure [64], by taking into account possible nontrivial endomorphisms of the induced module that dictates how such an induced module splits into simple $A$-modules. [^25] This is the direct generalization of the usual fixed point resolution in ordinary orbifold [103] for both local operators and defects. Here this splitting happens for the two rank 4 module categories, where the induced module $A\otimes{\mathcal{V}}$ decomposes into two simple $A$-modules that are differentiated by the subscript on $\mathcal{V}$}, and the corresponding NIM-irreps are given in (3.12) and (3.13). For the algebra object $A=\mathds{1}\oplus\eta\oplus\eta^{\prime}\oplus\eta\eta^{\prime}$ there are two inequivalent multiplication morphisms that differ by the discrete torsion $H^{2}({\mathbb{Z}}_{2}^{2},U(1))={\mathbb{Z}}_{2}$ (differentiated by the subscript $\star$ in Table 1). One of them is Morita equivalent to trivial gauging where the multiplication morphism is the trivial canonical one for $A={\mathcal{V}}\otimes{\mathcal{V}}$, and the induced module $A\otimes{\mathcal{V}}$ splits into four simple $A$-modules. The resulting module category is the regular module category that is universal to any fusion category with the corresponding NIM-irrep (3.11). In the other case, this splitting is forbidden due to the nontrivial multiplication morphism, and the resulting module category is rank 2 and corresponds to the NIM-irrep (3.15). Finally the two non-invertible gaugings where ${\mathcal{V}}\in A$ produce a rank 2 module category and a rank 1 module category with NIM-irreps (3.14) and (3.16), respectively. The rank 1 module category $\rm\bf Vec$} with the unique maximal algebra object realizes the unique fiber functor for ${\rm\bf Rep}(H_{8})$ [90].
@@
where the ${\rm\bf Rep}(H_{8})$-module category ${\mathcal{M}}(\hat{\mathbb{Z}}_{2},1)$ is equivalent to the opposite of the ${\rm\bf Vec}_{D_{8}}^{\gamma}$-module category ${\mathcal{M}}({\mathbb{Z}}_{2}^{s},1)$ as an invertible bimodule category over $({\rm\bf Rep}(H_{8}),{\rm\bf Vec}_{D_{8}}^{\gamma})$. All ${\mathbb{Z}}_{2}$ symmetries in ${\rm\bf Rep}(H_{8})$ are clearly gaugeable, however the special one $\hat{\mathbb{Z}}_{2}$ needed here must be the subgroup generated by the invertible TDL $\eta\eta^{\prime}$ which is distinguished by having trivial F-symbols (i.e. 1 if the fusion channel exists) with all TDLs in ${\rm\bf Rep}(H_{8})$. This ensures that the duality defect $\mathcal{V}$} under this $\hat{\mathbb{Z}}_{2}$ gauging can split consistently into a pair of conjugate invertible TDLs that correspond to generators of the ${\mathbb{Z}}_{4}$ subgroup of dual $D_{8}$ symmetry. We thus identify ${\rm\bf Vec}_{D_{8}}^{\gamma}$ as the dual category for gauging $A=\mathds{1}\oplus\eta\eta^{\prime}$ in Table 1.
@@
as is found in [105]. The full groupoid $\rm\bf BrPic$} consists of $8\times 2^{2}=32$ invertible bimodule categories (Morita equivalence relations) whose composition laws are all commutative.
@@ 3.4 Algebra objects in ${\rm\bf Rep}(D_{8})$
$L\subset D_{8}$ Algebra object AA Module Category ${\mathcal{C}}_{A}$ Dual Fusion Category $_{A}{\mathcal{C}}_{A}$
@@
$\langle s\rangle$ $\mathds{1}\oplus\eta\eta^{\prime}\oplus{\mathcal{V}}$ ${\rm NR}_{4}$: $\{A,\eta\oplus\eta^{\prime}\oplus{\mathcal{V}}\}$ ${\rm\bf Rep}(D_{8})$
$\langle sr\rangle$ $\mathds{1}\oplus\eta^{\prime}\oplus{\mathcal{V}}$ ${\rm NR}_{4}$: $\{A,\eta\oplus\eta\eta^{\prime}\oplus{\mathcal{V}}\}$ ${\rm\bf Rep}(D_{8})$
@@
Table 2: Algebra objects $A$ in ${\rm\bf Rep}(D_{8})$ grouped according to the Morita equivalence classes together with the corresponding module categories ${\mathcal{C}}_{A}$ (simple objects therein and the corresponding NIM-rep from Section 3.1) and dual fusion categories $_{A}{\mathcal{C}}_{A}$. The subscripts on the algebra object AA label different multiplication morphisms. The first column labels the module categories using group-theoretical data in $D_{8}$ according to (2.28) for $(H,\psi)=(D_{8},1)$.